The probabiliy of meetng srandom prowh has the ame bi (1 point) 365 is approximately 0.27%. What is the probability that the 25th person you meet is the first person who has the same birthday as you? 0.27% 00.14% 00.25% 0.03% 2, Keyshawn is bowling in a competition. He has a 26% chance of getting a strike each time he (I point) bowls. What is the probability that Keyshawn gets a strike in his first 5 attempts? 00.22 00.25 00.78 00.81 3. A computer programmer has an 18% chance of finding a bug in any given program. What is the probability that she does not find a bug within the first 10 programs she examines? (Ipino 00.03 00.14 0.44 00.86 4. Donnell is doing market research for a major brand by surveying customers at a local mall. ( point) The probability of a customer agreeing to take the survey is 22%, what is the probability that one of the first five customers agrees to take the survey? 0.29% 054% 071% C, 74%
Answer #1

1)
This is geometric distribution
P(X = k) = (1-p)^(k-1) * p
here
k = 25 , p= 0.0027
hence
P(X = 25) = (1-0.0027)^24 * 0.0027
= 0.00253036
=0.25 %

option C) is correct

2)
1 - P(X = 0)
= 1- (1-p)^n
= 1 - (1 - 0.26)^5
= 0.7780993
option C) 0.78

3)
(1-p)^n
= (1-0.18)^10
= 0.1374480313

B) 0.14

4)
1- (1-p)^n
= 1- (1-0.22)^5
= 0.71128
C) 71%

Please rate